# Pathspace Kalman Filters with Dynamic Process Uncertainty for Analyzing Time-course Data

### v1

@@ Abstract
Kalman Filter (KF) is an optimal linear state prediction algorithm, with applications in fields as diverse as engineering, economics, robotics, and space exploration. Here, we develop an extension of the KF, called a Pathspace Kalman Filter (PKF) which allows us to a) dynamically track the uncertainties associated with the underlying data and prior knowledge, and b) take as input an entire trajectory and an underlying mechanistic model, and using a Bayesian methodology quantify the different sources of uncertainty. An application of this algorithm is to automatically detect temporal windows where the internal mechanistic model deviates from the data in a time-dependent manner. First, we provide theorems characterizing the convergence of the PKF algorithm. Then, we numerically demonstrate that the PKF outperforms conventional KF methods on a synthetic dataset lowering the mean-squared-error by several orders of magnitude. Finally, we apply this method to biological time-course dataset involving over 1.8 million gene expression measurements.
@@ 1 Introduction
Estimating sources of uncertainty about a system from data is a quintessential problem with applications in many branches of science and engineering [9], [42]. The Kalman Filter (KF) is a highly successful and provably optimal solution to this problem under certain assumptions [1]. Crucially, KFs carefully account for different sources of noise, including data variance, model variance, and model confidence (called process uncertainty) in order to yield optimal estimates. Figure 1A illustrates the key components of a KF, and Figure 1B illustrates the dynamic state prediction of a standard KF. The KF formalism is quite flexible and allows diverse domain-specific internal models coupled with a general numerically tractable estimation procedure. This flexibility of KFs has allowed for its successful application in diverse fields such as weather prediction [16], target tracking [17], data smoothing and assimilation [6], flight [25], parameter estimation in electro-chemical responses [3], microscopy [18], image processing [46], economics [4], music retrieval [29], and more.
@@
More generally, KFs can be seen as simple probabilistic graphical models [14]. In this light, the KF models a system as a normal distribution with mean and variance derived as a convex combination of a measurement normal distribution and the internal model normal distribution, (Figure 1A) allowing the relative contributions of noise from these sources to be quantified. Kalman Filters are traditionally used to iteratively predict the state of a system at a future time given a stream of incoming measurements (Figure 1B). In many time-course experiments where KFs are used as an analysis tool, the entire measurement trajectory may be collected prior to analysis [22], [26, 30], [11, 20].
However, most methods fail to make use of the entire data trajectory simultaneously when optimizing the KF and instead iteratively run the filter along the data forward in time. Any procedure that fails to capitalize on the availability of the entire time trajectory may result in less than optimal estimates if future data may somehow constrain earlier predictions. Moreover, in a context with dynamical data and a potentially time-varying generative process, as is common in biology, there should be no apriori assumptions about the confidence in the underlying model being equal at all times times. Rather, we propose estimating the process uncertainty at different times in order to dynamically quantify the reliability of the internal model. For example, if there is a change in the underlying data generation process which is not accounted for in the internal model, the process uncertainty should adapt to this discrepancy. These observations motivate the development of a KF which uses information from an entire trajectory iteratively (Figure (1C)) to improve its estimates, including dynamically estimating process uncertainty, as opposed to only iterating forward in time.
@@
1. Pathspace Kalman Filter (PKF): In Section 3.1, we describe the Pathspace Kalman Filter, which analyzes the entire time-series at once and dynamically quantifies process uncertainty. We prove the PKF allows for non-monotonic process uncertainty, yet that the PKF algorithm converges.
2. Bayesian Computation: We describe a scalable Bayesian methodology for simulation based computation of model expectation and variance in Section 4. This allows the PKF to be run efficiently on large datasets such as transcriptomics data.
3. Synthetic Data Experiments: In Section 5 we compare a standard KF and our PKF models on synthetic data and show a decrease in mean-squared-error by several orders of magnitude.
@@
Figure 1: A. Illustration of the Kalman filtering framework. A mechanistic model produces a distribution of predictions which are combined with the underlying data and process uncertainty to produce a low variance state estimate (filter prediction). B. Illustration of a standard Kalman filter which makes predictions forward in time as new data is acquired. C. Illustration of a pathspace Kalman filter which ingests the entire temporal trajectory to produce a an entire trajectory as output.
@@ 2 Background
2.1 Kalman Filters
A 1-dimensional Kalman Filter (KF) is an optimal prediction-measurement-update algorithm in the sense that given a linear system with added Gaussian noise, the KF will produce the lowest variance estimate from the underlying data [1], [43]. Internally, a KF achieves this by linearly combining the predictions of an internal model $m$ with the observed data in order to fit a minimum variance normal distribution. Let $Z_{t}=Y_{t}+N_{t}$ be a sequence of data and $M_{t}\sim m(Z_{t-1})$ be a corresponding sequence of model predictions with well defined means and variances. Here, $Y_{t}$ are the true values, $N_{t}$ are i.i.d. Gaussian noise, and $t$ indexing time-points, and $m$ is the internal model of the KF. Then the KF prediction is given by:
$w_{t}=\underset{w^{\prime}}{\textrm{argmin}}[w_{t}^{\prime 2}\operatorname{\sigma^{2}}(Z_{t-1})$
$+(1-w_{t}^{\prime})^{2}(\operatorname{\sigma^{2}}(M_{t})+Q)]$ (3)
where, the terms are as described in 1. $\operatorname{\mathbb{E}}(F_{t})$ denotes the filter estimate of $Y_{t}$, which, in standard KFs is the minimum variance estimate of the process $Y_{t}$. The process uncertainty, $Q$, is traditionally used as a parameter to bias the relative weight between model and observations. Finally the weight $w_{t}$ is the Kalman gain which is monotonically decreasing in $t$ provided $Q$ and $\operatorname{\sigma^{2}}(Z_{t})$ are constant [1]. In the case of a linear model $m$, this monotonicity gives rise to the convergence of the filter variance:
The Kalman gain represents new information gained with each subsequent measurement [43]. A standard KF is depicted as a graphical model in Figure (2A).
@@ 2.1 Kalman Filters
STANDARD KALMAN FILTER (KF)
@@
$\operatorname{\boldsymbol{m}}$ Internal model of the PKF
@@
3 Preliminaries
3.1 Pathspace Kalman Filter (PKF)
To motivate the PKF, we first note that the KF only looks forward in time. In many applications, the entire time-series dataset is already available, and it is desirable that all the dynamics are analyzed together. To do this, we define the PKF which can be seen as an adaptation of the KF where the entire filter iterates over its dynamical output (the path) repeatedly. Additionally, the PKF dynamically updates process uncertainty in order to re-weight the model confidence. Specifically the update equations are:
$\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}_{t}^{i})=\operatorname{\boldsymbol{w}}^{i}_{t}\operatorname{\mathbb{E}}(Z_{t})+\operatorname{\boldsymbol{v}}_{t}^{i}\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{M}}^{i}_{t})+\operatorname{\boldsymbol{u}}_{t}^{i}\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}^{i-1}_{t}),$ (5)
$\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}_{t}^{i})=(\operatorname{\boldsymbol{w}}_{t}^{i})^{2}\operatorname{\sigma^{2}}(Z_{t})$
$+(\operatorname{\boldsymbol{v}}_{t}^{i})^{2}(\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})+\operatorname{\boldsymbol{Q}}_{t}^{i-1})+(\operatorname{\boldsymbol{u}}_{t}^{i})^{2}\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}^{i-1}_{t}),$ (6)
@@ 3.1 Pathspace Kalman Filter (PKF)
$\operatorname{\boldsymbol{w}}^{i}_{t},\operatorname{\boldsymbol{v}}^{i}_{t}=\underset{\operatorname{\boldsymbol{w}}^{\prime i}_{t},\operatorname{\boldsymbol{v}}^{\prime i}_{t}}{\textrm{argmin}}[(\operatorname{\boldsymbol{w}}^{\prime i}_{t})^{2}\operatorname{\sigma^{2}}(Z_{t})$
$+(\operatorname{\boldsymbol{v}}^{\prime i}_{t})^{2}(\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})+\operatorname{\boldsymbol{Q}}_{t}^{i-1})+(\operatorname{\boldsymbol{u}}^{\prime i}_{t})^{2}\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}^{i-1}_{t})],$
$\operatorname{\boldsymbol{u}}^{i}_{t}=1-\operatorname{\boldsymbol{w}}_{t}^{i}-\operatorname{\boldsymbol{v}}_{t}^{i},$ (7)
@@
There are three different weights $\operatorname{\boldsymbol{w}}$, $\operatorname{\boldsymbol{v}}$ and $\operatorname{\boldsymbol{u}}$ for the data, model, and filter, respectively. The symbols have been made bold to denote that they are vectors in both time $t$ (subscript) and filter iteration $i$ (superscript). For example $\operatorname{\boldsymbol{F}}^{i}=(\operatorname{\boldsymbol{F}}^{i}_{0},...,\operatorname{\boldsymbol{F}}^{i}_{T})$ with $T$ the number of time-points. Notice that the model samples $\operatorname{\boldsymbol{M}}^{i}_{t}\sim\operatorname{\boldsymbol{m}}(\operatorname{\boldsymbol{F}}^{i-1},t)$ are now a function of the same time $t$ and the entire trajectory $\operatorname{\boldsymbol{F}}^{i-1}$ from the previous filter iteration instead of the single filter value $F_{t-1}$ at the previous time. Similarly, the process uncertainty $\operatorname{\boldsymbol{Q}}_{t}^{i-1}$ is now a function of time and iteration instead of a constant. instead of $Q_{t-1}$. Feeding the filter output back into itself is important to ensure the convergence of the PKF as we prove in Appendix 8.1.2. The PKF is illustrated as a graphical model in Figure (2B).
@@
This entire construction can be viewed as an instantiation of the expectation maximization algorithm where in expectation step $i-1$ the filter output $\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}^{i-1})$ is computed. This value is then fed into the maximization step $i$ via the model $\operatorname{\boldsymbol{M}}^{i}\sim\operatorname{\boldsymbol{m}}(\operatorname{\boldsymbol{F}}^{i-1})$ where the filter variance is minimized to compute $\operatorname{\boldsymbol{w}}_{t}^{i}$. The new Kalman gain can then be used to compute the filter output $\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}_{t}^{i})$ and so on.
@@
In the KF formalism, the Kalman gain $\operatorname{\boldsymbol{w}}_{t}$ can be interpreted as the amount of new information gained from the most recent measurement. This value decreases monotonically for a standard KF, meaning that the weighting of the uncertainty associated with the data can never increase in time [43]. Monotonicity of the Kalman gain in time is undesirable because it does not allow for dynamically varying (both increasing and decreasing) the relative contributions to uncertainties from the data and model over time. Next, we prove that for the PKF the Kalman Gain $\operatorname{\boldsymbol{w}}^{i}_{t}$ is monotonically decreasing in $i$ but not $t$.
@@
Proof Outline: [Detailed proof in Appendix 8.1.1]. For a given filter iteration, $i$, the Kalman Gain ($\operatorname{\boldsymbol{w}}^{i}_{t}$) is given by:
@@
where, $A=\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}^{i-1}_{t})$, $B=(\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})+\operatorname{\boldsymbol{Q}}^{i-1})$, and $C=\operatorname{\sigma^{2}}(Z)$. As $\operatorname{\boldsymbol{Q}}_{t}^{i-1}$ and $\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}_{t}^{i})$ are not restricted to be monotonic in time, it follows that the sequence of weights, $\operatorname{\boldsymbol{w}}^{i}_{1}$, $\operatorname{\boldsymbol{w}}^{i}_{2}$,…, $\operatorname{\boldsymbol{w}}^{i}_{t}$} may be non-monotonic in time.
@@
Proof Outline: [Detailed proof in Appendix 8.1.2]. The non-monotonicity of $\operatorname{\boldsymbol{Q}}^{i}_{t}$ originates in the non-monotonicity of the loss function $L$ referenced in Equation (8) when applied to dynamical data. The fact that the process uncertainty can vary over the course of a trajectory allows $\operatorname{\boldsymbol{Q}}^{i}_{t}$ to be used as a dynamic metric to assess the relative sources of uncertainty from the model and data.
@@
Proof Outline: [Detailed proof in Appendix 8.1.2]. First, we show that the filter variance is monotonically decreasing towards its fixed point $\lim_{i\to\infty}\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}_{t}^{i})$. Then, we show the existence of a fixed point for $\lim_{i\to\infty}\operatorname{\boldsymbol{Q}}_{t}^{i}$ and show that convergence of the filter variance guarantees that the process uncertainty also converges.
Figure 2: Graphical model illustrations of Kalman Filter architectures. Solid arrows indicate dependencies of means and variances. Dashed arrows indicate dependencies only of the variance. A. Graphical Model of a standard Kalman Filter (KF). B. Graphical Model of the path-space Kalman Filter (PKF). Notice that the PKF feeds into itself using the index $i$ instead of in $t$ allowing for an entire trajectory to be iteratively processed.

### v2

@@ Abstract
Kalman Filter (KF) is an optimal linear state prediction algorithm, with applications in fields as diverse as engineering, economics, robotics, and space exploration. Here, we develop an extension to a non-linear adaptive KF closely related to Bayesian smoothing algorithms, called a Pathspace Kalman Filter (PKF) which allows us to a) dynamically track and update the uncertainties associated with the underlying data and prior knowledge, and b) take as input an entire trajectory and an underlying mechanistic model, and using a Bayesian methodology quantify the different sources of uncertainty. An application of this algorithm is to automatically detect temporal windows where the internal mechanistic model deviates from the data in a time-dependent manner. First, we provide theorems characterizing the convergence of the PKF algorithm. Then, we numerically demonstrate that the PKF outperforms state-of-the-art methods on a synthetic dataset by over an order of magnitude. Finally, we apply this method to biological time-course dataset involving over 1.8 million gene expression measurements.
@@ 1 Introduction
Estimating sources of uncertainty about a system from data is a quintessential problem with applications in many branches of science and engineering [rao], [kf_hep]. The Kalman Filter (KF) is a provably optimal solution to this problem under certain assumptions [kalman]. Crucially, KFs carefully account for different sources of noise, including data variance, model variance, and model confidence (called process uncertainty) in order to yield optimal estimates. Figure 1A illustrates the key components of a KF, and Figure 1B and 1D illustrate the dynamic state prediction of a standard KF.
The KF formalism has been extended in many ways including adaptive [brown1985adaptive] and non-linear KFs [extendedkf, unscentedkf] which allow diverse domain-specific internal models coupled with a general numerically tractable estimation procedure. This flexibility of KFs has allowed for its successful application in diverse fields such as economics [harvey1990forecasting], weather prediction [galanis2006applications], target tracking [weng2006video], data smoothing and assimilation [NIPS1996_147702db], flight [5466132], parameter estimation in electro-chemical responses [BrownandRutan], microscopy [yang2006nuclei], image processing [nahata2023exploring], music retrieval [7042773], and more.
@@
More generally, KFs can be seen as simple probabilistic graphical models [jordan2004introduction]. In this light, the KF models a system as a normal distribution with mean and variance derived as a convex combination of a measurement normal distribution and the internal model normal distribution, (Figure 1A) allowing the relative contributions of noise from these sources to be quantified. Kalman Filters are traditionally used to iteratively predict the state of a system at a future time given a stream of incoming measurements (Figure 1B, 1E).
In many time-course experiments where KFs are used as an analysis tool, the entire measurement trajectory may be collected prior to analysis [Jalles2009], [Lillacci2010, Huang2016], [Wu2002, Eden2008]. In such cases, Bayesian smoothing algorithms may improve state estimation by utilizing the entire trajectory. These algorithms frequently use KFs internally to iterate over the entire data trajectory and generate dynamic internal estimates of sources of uncertainty [sarkka2023bayesian]. However, such approaches are sensitive to algorithm parameters which frequently must be predetermined by the user.
In the context of modeling biological phenomena as time-varying generative process, there should be no apriori assumptions about the confidence in the underlying model being equal at all times times. In the KF formalism, this confidence is codified in the process uncertainty. In this work, we propose adaptively estimating the process uncertainty based on a loss function in order to dynamically quantify the reliability of the internal model. For example, if there is a change in the underlying data generation process which is not accounted for in the internal model, the process uncertainty should adapt to this discrepancy. These observations motivate the development of an algorithm which uses information from an entire trajectory (Figure (1C, 11D)) to improve its estimates. Our approach is to define a KF in pathspace which feeds trajectory estimates back into itself iteratively. We call our algorithm a Pathspace Kalman Filter (PKF), because, unlike a Bayesian smoother, our goal is not to smooth the data but rather to minimize state estimation variance and error across the entire trajectory. Finally, we note that an additional advantage of our approach is that it automatically allows for change point detection - meaning identifying when changes in a data generation process occur [aminikhanghahi2017survey].
@@
1. Pathspace Kalman Filter (PKF): In Section 3, we describe the Pathspace Kalman Filter, which is a univariate non-linear adaptive KF that analyzes the entire time-series and dynamically quantifies process uncertainty while feeding its output trajectory (or path) back into itself, which we call iterating in pathspace. We prove the PKF allows for non-monotonic process uncertainty in time, yet that the PKF algorithm converges as it iterates in pathspace.
2. We describe a scalable Bayesian methodology for simulation based computation of model expectation and variance in Section 4. This allows the PKF to be run efficiently on large datasets such as transcriptomics data.
3. Synthetic Data Experiments: In Section 5 we compare several KF algorithms and Bayesian smoothing algorithms with the PKF models on synthetic data and show that the PKF outperforms all other algorithms by one or more orders of magnitude in terms of mean-squared-error and requires less hyper parameter optimization.
@@
Figure 1: A. A schematic of the Kalman filtering framework. B. Graphical model for the KF. C. Graphical model of the proposed PKF. D. Illustration of a KF which makes predictions forward in time $t$ as new data is acquired. E. Illustration of a PKF which ingests the entire temporal trajectory, over all $t$, to produce an entire trajectory as output. At each iteration $i$, the PKF output trajectory is fed back into the PKF as the input for the next iteration $i+1$ of filtering.
@@ 2 Background
ADAPTIVE NON-LINEAR KALMAN FILTER (KF)
@@
$\operatorname{\boldsymbol{m}}(\operatorname{\boldsymbol{F}}^{i},t)$ Internal model of the PKF
@@
2.1 Kalman Filters
A 1-dimensional Kalman Filter (KF) is an optimal prediction-measurement-update algorithm in the sense that given a linear system with added Gaussian noise, the KF will produce the lowest variance estimate from the underlying data [kalman], [kalmanfilternet]. Internally, a KF achieves this by linearly combining the predictions of an internal model $m$ with the observed data in order to fit a minimum variance normal distribution. In this work, we are interested in non-linear univariate KFs, so we do not require $m$ to be linear. Let $Z_{t}=Y_{t}+N_{t}$ be a sequence of data and $M_{t}\sim m(Z_{t-1})$ be a corresponding sequence of samples of model predictions with well defined means and variances. Here, $Y_{t}$ are the true values, $N_{t}$ are i.i.d. Gaussian noise, and $t$ indexing time-points, and $m$ is the internal model of the KF. Then the KF prediction is given by:
$w_{t}$ $=\underset{w^{\prime}}{\textrm{argmin}}[w_{t}^{\prime 2}\operatorname{\sigma^{2}}(Z_{t-1})+(1-w_{t}^{\prime})^{2}(\operatorname{\sigma^{2}}(M_{t})+Q)]$ (3)
where, the terms are as described in Table 1. $\operatorname{\mathbb{E}}(F_{t})$ denotes the filter estimate of $Y_{t}$, which, in standard KFs is the minimum variance estimate of the process $Y_{t}$. The process uncertainty, $Q$, is traditionally used as a parameter to bias the relative weight between model and observations. Finally the weight $w_{t}$ is the Kalman gain which is monotonically decreasing in $t$ provided $Q$ and $\operatorname{\sigma^{2}}(Z_{t})$ are constant [kalman]. In the case of a linear model $m$, this monotonicity gives rise to the convergence of the filter variance:
The Kalman gain represents new information gained with each subsequent measurement [kalmanfilternet]. Note that in the original KF, the model $m$ is assumed to be a linear function. Extensions of this model such as the extended KF and unscented KF relax this constraint and allow for non-linear internal models [extendedkf, unscentedkf]. In this text, we will generally use KF to mean a non-linear adaptive Kalman filter so that $m$ is not required to be linear and that the data variance $\operatorname{\sigma^{2}}(Z)$ is dynamically updated.
3 Pathspace Kalman Filter (PKF)
To motivate the PKF, we recognize that biological time-series analyses would benefit from an algorithm that is able to use an entire trajectory, looking forwards and backwards in time, to compute the process uncertainty dynamically. Such an algorithm should be able to highlight temporal windows where the data generative process deviates from our internal model. We define the PKF as an extension of a non-linear adaptive univariate KF where the entire filter iterates over its dynamical output (the path) repeatedly in a procedure similar to iterative Bayesian smoothing approaches such as the Iterated Posterior Linearization Smoothing (IPLS) algorithm [IPLS]. Additionally, the PKF dynamically updates process uncertainty in order to re-weight the model confidence.
Our update equations are similar to a KF’s with two conceptual changes. First, the filter’s $i^{\textrm{th}}$ estimated trajectory distribution at time $t$, $\operatorname{\boldsymbol{F}}^{i}_{t}$ (equation (5)), is a convex combination of three Gaussian distributions instead of two: the data $(Z_{t})$, the internal model prediction $(\operatorname{\boldsymbol{M}}^{i}_{t})$, and the previous filter update $(\operatorname{\boldsymbol{F}}^{i-1}_{t})$, which allows the filter to feed back on itself. Second, the PKF dynamically updates process uncertainty $(\operatorname{\boldsymbol{Q}}^{i}_{t})$ using a loss function $L$ (such as the $L_{2}$ norm) between the model output and the underlying data in order to minimize the loss (equation (8)). Specifically the update equations are:
$\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}_{t}^{i})$ $=\operatorname{\boldsymbol{w}}^{i}_{t}\operatorname{\mathbb{E}}(Z_{t})+\operatorname{\boldsymbol{v}}_{t}^{i}\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{M}}^{i}_{t})+\operatorname{\boldsymbol{u}}_{t}^{i}\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}^{i-1}_{t})$ (5)
$\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}_{t}^{i})$ $=(\operatorname{\boldsymbol{w}}_{t}^{i})^{2}\operatorname{\sigma^{2}}(Z_{t})+(\operatorname{\boldsymbol{v}}_{t}^{i})^{2}(\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})+\operatorname{\boldsymbol{Q}}_{t}^{i-1})+(\operatorname{\boldsymbol{u}}_{t}^{i})^{2}\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}^{i-1}_{t})$ (6)
@@ 3 Pathspace Kalman Filter (PKF)
$\operatorname{\boldsymbol{w}}^{i}_{t},\operatorname{\boldsymbol{v}}^{i}_{t}$ $=\underset{\operatorname{\boldsymbol{w}}^{\prime i}_{t},\operatorname{\boldsymbol{v}}^{\prime i}_{t}}{\textrm{argmin}}\left[(\operatorname{\boldsymbol{w}}^{\prime i}_{t})^{2}\operatorname{\sigma^{2}}(Z_{t})+(\operatorname{\boldsymbol{u}}^{\prime i}_{t})^{2}\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}^{i-1}_{t})+(\operatorname{\boldsymbol{v}}^{\prime i}_{t})^{2}(\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})+\operatorname{\boldsymbol{Q}}_{t}^{i-1})\right]$
[…]
@@
There are three different weights $\operatorname{\boldsymbol{w}}$, $\operatorname{\boldsymbol{v}}$ and $\operatorname{\boldsymbol{u}}$ for the data, model, and filter, respectively. The symbols have been made bold to denote that they are vectors in both time $t$ (subscript) and filter iteration $i$ (superscript). For example $\operatorname{\boldsymbol{F}}^{i}=(\operatorname{\boldsymbol{F}}^{i}_{0},...,\operatorname{\boldsymbol{F}}^{i}_{T})$ with $T$ the number of time-points. Notice that the model samples $\operatorname{\boldsymbol{M}}^{i}_{t}\sim\operatorname{\boldsymbol{m}}(\operatorname{\boldsymbol{F}}^{i-1},t)$ are now a function of the same time $t$ and the entire trajectory $\operatorname{\boldsymbol{F}}^{i-1}$ from the previous filter iteration instead of the single filter value $F_{t-1}$ at the previous time. Similarly, the process uncertainty $\operatorname{\boldsymbol{Q}}_{t}^{i-1}$ is now a function of time and iteration instead of a constant. The prefactor $\operatorname{\boldsymbol{w}}^{i}_{t}+\operatorname{\boldsymbol{v}}^{i}_{t}$ in equation (8) is analogous to a Kalman Gain for the process uncertainty and weights its update proportional to the additional uncertainty the model requires in order to match the data.
Algorithm 1 Pathspace Kalman Filter
  Inputs: data $Z$, model $\operatorname{\boldsymbol{m}}$, iterations $I_{tot}$, loss function $L$, and total time points $T$
  Outputs: $\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}^{I_{tot}})$, $\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}^{I_{tot}})$, and $\operatorname{\boldsymbol{Q}}^{I_{tot}}$
  $\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}_{t}^{0})\leftarrow\operatorname{\mathbb{E}}(Z_{t})$Initialization
  $\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}_{t}^{0})\leftarrow\operatorname{\sigma^{2}}(Z_{t})$
  $\operatorname{\boldsymbol{Q}}_{t}^{0}\leftarrow\operatorname{\sigma^{2}}(Z_{t})$
  for iteration $i=1$ to $I_{tot}$  do
     for timepoint $t=0$ to $T$  do
        $\operatorname{\boldsymbol{M}}^{i}_{t}\leftarrow\operatorname{\boldsymbol{m}}(\operatorname{\boldsymbol{F}}^{i-1}_{t})$Model Prediction
        Compute $\operatorname{\boldsymbol{w}}_{t}^{i}$, $\operatorname{\boldsymbol{v}}_{t}^{i}$, $\operatorname{\boldsymbol{u}}_{t}^{i}$ with Eq. (3)
        Compute $\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}_{t}^{i})$ with Eq. (5) #Filter Update
        Compute $\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}_{t}^{i})$ with Eq. (6)
        Compute $L(\operatorname{\boldsymbol{M}}_{i}^{t},Z_{t})$Uncertainty Update
        Compute $\operatorname{\boldsymbol{Q}}_{t}^{i}$ via Eq. (8)
     end for
  end for
@@
Feeding the filter output back into itself is important to ensure the convergence of the PKF as we prove in Appendix A.1. The PKF is illustrated as a graphical model in Figure (1C). Additionally, although it appears complex, the solution to equation (3) can be solved analytically as shown in equations (24-26) in appendix A.1.2. This entire construction can be viewed as an instantiation of the expectation maximization algorithm where in expectation step $i-1$ the filter output $\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}^{i-1})$ is computed. This value is then fed into the maximization step $i$ via the model $\operatorname{\boldsymbol{M}}^{i}\sim\operatorname{\boldsymbol{m}}(\operatorname{\boldsymbol{F}}^{i-1})$ where the filter variance is minimized to compute $\operatorname{\boldsymbol{w}}_{t}^{i}$, $\operatorname{\boldsymbol{v}}_{t}^{i}$ and $\operatorname{\boldsymbol{u}}_{t}^{i}$. The new Kalman gain can then be used to compute the filter output $\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}_{t}^{i})$ and so on.
@@
In the KF formalism, the Kalman gain $\operatorname{\boldsymbol{w}}_{t}$ can be interpreted as the amount of new information gained from the most recent measurement. This value decreases monotonically for a standard KF, meaning that the weighting of the uncertainty associated with the data can never increase in time [kalmanfilternet]. Monotonicity of the Kalman gain in time is undesirable because it does not allow for dynamically varying (both increasing and decreasing) the relative contributions to uncertainties from the data and model over time. Next, we prove that for the PKF the Kalman Gain $\operatorname{\boldsymbol{w}}^{i}_{t}$ is monotonically decreasing in $i$ but not $t$.
@@
Proof Outline: [Detailed proof in Appendix A.1.1]. For a given filter iteration, $i$, the Kalman Gain ($\operatorname{\boldsymbol{w}}^{i}_{t}$) is given by:
@@
where, $A=\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}^{i-1}_{t})$, $B=(\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})+\operatorname{\boldsymbol{Q}}^{i-1})$, and $C=\operatorname{\sigma^{2}}(Z_{t})$. As $\operatorname{\boldsymbol{Q}}_{t}^{i-1}$ and $\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}_{t}^{i})$ are not restricted to be monotonic in time, it follows that the sequence of weights, $\operatorname{\boldsymbol{w}}^{i}_{1}$, $\operatorname{\boldsymbol{w}}^{i}_{2}$,…, $\operatorname{\boldsymbol{w}}^{i}_{t}$} may be non-monotonic in time.
@@
Proof Outline: [Detailed proof in Appendix A.1.2]. The non-monotonicity of $\operatorname{\boldsymbol{Q}}^{i}_{t}$ originates in the non-monotonicity of the loss function $L$ referenced in Equation (8) when applied to dynamical data. The fact that the process uncertainty can vary over the course of a trajectory allows $\operatorname{\boldsymbol{Q}}^{i}_{t}$ to be used as a dynamic metric to assess the relative sources of uncertainty from the model and data. In fact, as is apparent in equation (8), $\operatorname{\boldsymbol{Q}}^{i}_{t}=L(\operatorname{\boldsymbol{M}}_{i}^{t},Z_{t})$ is a fixed point of the PKF which quantifies a different error at each time-point.
@@
Proof Outline: [Detailed proof in Appendix A.1.2]. First, we show that the filter variance is monotonically decreasing towards its fixed point $\lim_{i\to\infty}\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}_{t}^{i})$. Then, we show the existence of a fixed point for $\lim_{i\to\infty}\operatorname{\boldsymbol{Q}}_{t}^{i}$ and show that convergence of the filter variance guarantees that the process uncertainty also converges. Numerical results showing the convergence of the PKF can be found in Appendix A.2.
An important feature of classical KFs is that in the linear case they minimize the means-squared error (MSE). In the next theorem, we show an analogous result for a class of PKFs that we call ”accurate linear PKFs”. Accurate linear PKFs have an internal model $\operatorname{\boldsymbol{m}}$ that is a linear function which is identical to the underlying process dynamics. A formal definition is provided in Appendix (A.1.3).
Theorem 3: An accurate linear PKF is optimal in the sense that it minimizes the MSE at each time-point $t$, and each iteration $i$.
Proof Outline: [Detailed proof in Appendix A.1.3]. First, we show that filter output can be written as the expected value of the true process $(Y_{t})$ with added Gaussian noise. Then, we prove that the accurate linear PKF update equations minimize the MSE.